I need a five step proof (statement and reasoning) to:
Given: 3y+5=x and x=y+1
Prove: x= -1 & y=-2
I'm working with definitions, theorems, and postulates. Thank you so much!
Answer by josgarithmetic(39799) 
 You might have more information showing how or why "five steps".
3y+5=(y+1), substitution principal
3y-y=1-5, additive inverse, combined steps
2y=-4, combined like-terms
y=-2, used multiplicative inverse property
x=-2+1, substitution into the other given equation
x=-1, number fact arithmetic computation
Six steps used.
|
|
|